Institut for Matematik og Datalogi Odense Universitet

{May 17, 1996 \atop JFB}

DM11 – F96 – Lecture 15

Lecture, May 15

We finished with algebra. (See the notes from lecture 13.)

Lecture, May 22

I will answer any questions on course material. If there are not many, I will probably briefly talk about graph coloring.